Lannister Manufacturing has a target debt-equity ratio of 0.52. Its cost of equity is 18 percent, and its cost of debt is 9 percent. If the tax rate is 33 percent, what is the company's WACC?
Answer #1

debt-equity ratio = 0.52

=>

debt/equity = 0.52

and

debt + equity = 1

=>

weight of equity = 25/38

weight of debt = 13/38

WACC = weight of equity * cost of equity + weight of debt * cost of debt * (1-tax rate)

= 25/38 * 18% + 13/38 * 9% * (1-0.33)

= 13.905%

= 13.91%